OpenStudy (anonymous):

write each interval as an inequality [-3,3]

OpenStudy (anonymous):

please help

OpenStudy (agreene):

So, generally brackets mean exclusive (cannot equal) and parenthesis mean can equal. In this case ]-3,3[ -3<x<3

OpenStudy (agreene):

if it were (-3,3) it would have been \[-3\le x \le 3\]

OpenStudy (anonymous):

but brackets forsure are \[\le \ge\]